# ActionReasoningBench: Reasoning about Actions with and without Ramification Constraints

### v2

@@ 1 Introduction
Table 1: Differences between ActionReasoningBench (Ours) and previous benchmarks on RAC. PlanBench (Valmeekam et al., 2024a) ; TRAC (He et al., 2023)
@@ 2 Related Works
Evaluating the reasoning capabilities of LLMs has become a key focus in NLP, with increasing efforts to create challenging benchmarks. Notable areas of interest include the evaluation of legal reasoning (Fei et al., 2023) (Guha et al., 2023), logical reasoning (Luo et al., 2024) (Han et al., 2024) (Patel et al., 2024) (Parmar et al., 2024) arithmetic reasoning (Cobbe et al., 2021) (Miao et al., 2021), temporal reasoning (Uddin et al., 2024) (Fatemi et al., 2024), and commonsense reasoning (Onoe et al., 2021) (Lin et al., 2021) (Geva et al., 2021) (Lourie et al., 2021). Despite this progress, RAC remains significantly under-explored, even though it plays a crucial role in several of these reasoning tasks, such as commonsense and legal reasoning.
RAC and Planning
Previous works, such as those by Valmeekam et al. (2024a), Valmeekam et al. (2024b), and Guan et al. (2023), have curated benchmarks assessing the performance of LLMs on planning, with limited exploration of RAC. Banerjee et al. (2020) examined the RAC capabilities of LLMs across four domains, focusing only on three question types. Similarly, He et al. (2023) tested LLMs on RAC in a modified version of Blocksworld, employing four question categories—two of which centered on RAC, while the other two emphasized planning. In contrast, our work explores 8 distinct domains and includes 6 different question categories. Additionally, we introduce novel ramifications constraints, adding complexity to the RAC task by introducing indirect effects of actions.
@@ 3 ActionReasoningBench
This section provides a detailed overview of our benchmark, including its categorization, creation methodology, and validation process. A sample instance is presented in Appendix A, where we also describe the objects, actions, and fluents within the domain.
@@ 3.1 Question Categories
    State Tracking - This category builds upon and extends the concept of Fluent Tracking. It involves querying about the complete set of fluents in the final state. For instance, in the Blocksworld domain, a state-tracking question might be “What are all the valid properties in this state?”
@@
    Action Executability - This category encompasses two types of questions related to determining the executability of actions. The types of questions within this category are as follows:
@@ 3.2 Fluent Categories
Classification of every fluent across all 8 domains can be found in Appendix H. Furthermore, for each fluent type, we generate questions involving negative fluents, which allows us to evaluate LLM’s comprehension of negation within RAC contexts.
@@ 3.3 Dataset Structure and Variations
The ActionReasoningBench requires domains that facilitate the evaluation of LLMs on both short and long sequences of meaningful actions, where the effects and preconditions of these actions are succinctly described. Additionally, these domains should reflect real-world scenarios. To meet these criteria, we selected 8 domains Blocksworld, Depots, Driverlog, Grippers, Mystery, Satellite, Spanner, and Visitall-sourced from the International Planning Competition (IPC), covering the years 1998 to 2014. These domains are commonly used as benchmarks for evaluating advanced planning systems and provide a robust foundation for research in automated planning. Appendix F provides a detailed description of each domain. Notably, even state-of-the-art LLMs like GPT-4o are not capable of generating diverse domains or action sequences that conform to the precise constraints outlined in these domain descriptions, which justifies the reliance on IPC domains.
@@
The domains provided by the IPC are described using the Planning Domain Definition Language (PDDL), a formal language designed to model deterministic actions and state transitions for planning problems. Further information on PDDL can be found in Appendix G.1. In this study, we initially translated the PDDL-based domains into natural language. Subsequently, ramification constraints were introduced into these natural language descriptions. The process underwent validation by two experts in the RAC domain to ensure correctness. Given that the category Action Executability focuses on determining whether the action can be performed rather than analyzing its effect, we concentrated on the categories Fluent Tracking, State Tracking, and Effects of Actions when introducing ramifications.
@@
We formulate two distinct types of questions based on the expected answer format. The first type consists of binary questions, where the response is constrained to either “True” or “False”. The second type involves subjective questions, where the answer may encompass a range of multiple objects, actions, or fluents.
@@ 3.4 Data Creation & Validation
The question generation process follows a four-stage pipeline, as illustrated in Figure 1. The selected domains from the IPC are represented in PDDL (see Appendix F for examples). First, these PDDL representations are used to generate 10 pairs of initial and goal conditions. A PDDL solver (Helmert, 2006) and validator (Howey et al., 2004) are then employed to obtain and validate the action sequences necessary to transition from the initial to the goal state. In the second stage, the PDDL domains, instances, and action sequences are converted into Answer Set Programming (ASP) descriptions using Python-based templates.
In the third stage, ASP solvers are used to generate the action-state space and extract fluents for each state, along with identifying all executable and inexecutable actions. Further details on these formal languages are provided in Appendix G. Finally, the fourth stage involves converting the action-state data into questions using a Python template. Up to three natural language variations are created for every object, action, and fluent. These templated sequences are then paraphrased to Llama-3.1-70B-Instruct to ensure they sound natural and avoid repetition in long action sequences. Three independent annotators review both the templated and the paraphrased versions to assess their naturalness, as detailed in Appendix C. Additionally, all eight domain descriptions were manually translated from PDDL to natural language.
@@ 3.5 Data Splits
Table 2: Overview of the test and train sets of ActionReasoningBench across Question and Fluent Categories. The “Mixed Fluents” category represents questions that involve more than one type of fluent.
@@ 4 Experiments and Evaluation
To evaluate our benchmark, we tested four different LLMs and employed two prompting techniques. The LLMs include two proprietary state-of-the-art models–GPT-4o (Achiam et al., 2023) and o1-preview (OpenAI, 2024)–alongside two open-source models, Llama-3.1-8B-Instruct and Llama-3.1-70B-Instruct (Dubey et al., 2024). Each LLM was evaluated using both few-shot prompting with three examples (few-shot-3) Brown (2020) and zero-shot-CoT Kojima et al. (2022) prompting. While the entire dataset requires reasoning abilities, the ramification subset involves the most complex and challenging reasoning tasks. Given that o1-preview is specifically optimized for reasoning tasks and incurs significantly higher costs compared to GPT-4o [^1] As of Oct 2024, o1-preview is six times more expensive than GPT-4o https://openai.com/api/pricing/, we restricted its use to the ramification subset, where its superior reasoning capabilities are expected to provide the greatest benefit. Utilizing o1-preview across the entire dataset would not be cost-effective, as its advantages would be less pronounced for simpler reasoning tasks.
@@
ActionReasoningBench includes two types of answer formats, as outlined in Section 3.3: binary (true/false) and free-form responses. The evaluation process was performed separately for each answer type. For binary questions, we extracted “true” and “false” keywords from the model’s response and compared them to the ground truth. Since free-form answers can’t be evaluated using exact string matching, we employed human evaluation for the ramification questions. While human evaluation is highly accurate, it is not scalable, so we used Llama-3.1-70B-Instruct to assess all free-form responses. The specific prompt used to evaluate the LLMs and the correlation between Llama-3.1-70B-Instruct and human evaluations are detailed in Appendix E. For all experiments, we report the accuracy along with the standard error of the mean (SEM), calculated as $SEM=\frac{\sigma}{\sqrt{n}}$, where $\sigma$ represents the standard deviation, and $n$ is the sample size.
@@
We fine-tuned the Llama-3.1-8B model using the training data split outlined in Section 3.5. Due to the limited computing power, we excluded any data samples that exceeded a context length of 4096 tokens. The fine-tuning process was performed separately for free-response and binary questions. Detailed information on the fine-tuning procedure is provided in Appendix D. All experiments were executed using 8×H100 GPUs.
@@ 5 Results and Discussion
This section presents the results and analysis using ActionReasoningBench as defined in Section 4. The Zero-shot-CoT results for each LLM on both the binary and free-response subsets of the test set are provided in Tables 3 and 4 respectively. Similarly, the Few-shot-3 results for each LLM evaluated on these same subsets are displayed in Tables 6 and 7. The detailed analysis of the effects when ramification constraints are incorporated into the descriptions is discussed in Section 5.1.
@@
In our evaluation, GPT-4o demonstrated the highest performance on the Grippers domain and the lowest on the Satellite domain, with a performance gap of 15.53%. For both Llama-3.1-8B-Instruct and Llama-3.1-70B-Instruct, the best performance is also on the Grippers domain, but their worst performance occurs on the Mystery domain, with differences of 19.24% and 24.04%, respectively. Interestingly, the second lowest performing domain for GPT-4o is Mystery, while for the Llama models, it is Satellite. This suggests that despite potential differences in pre-training data, these models exhibit a similar relative understanding of the domains.
Decrease in Performance on Complex Questions
The categories of Composite Questions and Numerical RAC build upon the four foundational RAC dimensions. The performance of GPT-4o, Llama-3.1-8B-Instruct, and Llama-3.1-70B-Instruct on free-response complex questions highlights significant challenges, especially in the Numerical RAC category. This category reformulates existing question types into numerical formats, a domain where all tested LLMs exhibit notable difficulty. Specifically, performance on numerical questions related to Action Executability is 8.65% lower than on questions in the Fluent Tracking category. Previous research, such as (Ahn et al., 2024), indicates that LLMs struggle with arithmetic reasoning, which, when mixed with the RAC questions, likely contributes to the poor performance in the Numerical RAC category.
@@
Across all LLMs examined in the study, a consistent trend emerges in which performance on Static Properties decreases as the length of action-sequence increases. While these static properties remain unchanged throughout the actions, they might get overlooked in longer sequences, likely due to their absence in the effect of any action. This phenomenon resembles the “needle in a haystack” challenge in long-context scenarios, where LLMs struggle to recall specific information embedded within a long context (Zhang et al., 2024). Conversely, Base Fluents maintain stable performance across all action sequences, indicating that the LLMs consistently capture the direct effects of actions. Ramification fluents exhibit a steady decline in performance as the sequence lengthens, particularly affecting the subcategory of Derived Fluents, which suggests that LLMs have more difficulty handling indirect effects. Finally, Mixed Fluents, which involve more than one fluent type, show a consistent decline in performance as the length of action sequences increases.
@@
Figure 2: Question Categories for both T/F and Free answers as a function of Action-Sequence length for 3 models (zero-shot-CoT). Note: bars represent SEM.
Figure 2 illustrates the performance of the three models–GPT-4o, Llama-3.1-70B-Instruct, and Llama-3.1-8B-Instruct–across varying action sequence lengths in a zero-shot-CoT setting. The results combine both binary and free-response formats, with detailed data available in Tables 3 and 4. Generally, model accuracy declines as the action-sequence length increases, a pattern that holds for most categories. However, GPT-4o deviates from this trend in the State Tracking category, where performance first decreases and then improves. Since this trend is not observed with the other two models, and the results lie within the margin of error, we believe this is an outlier. In contrast, the Effects of Actions category consistently deviates from this trend, likely due to the nature of the task, which focuses on changes resulting from the last action, making it less dependent on the sequence of actions.
@@
As demonstrated in Tables 3 and 4, the Llama-3.1-70B-Instruct model consistently outperforms the smaller Llama-3.1-8B-Instruct model, with an average performance improvement of 20.84%. This improvement is likely due to the larger model’s superior reasoning capabilities stemming from its increased number of parameters. A similar trend is observed when comparing GPT-4o to Llama-3.1-70B-Instruct, where GPT-4o exhibits an average performance increase of 8.64%. Although the specific size of GPT-4o remains undisclosed, it is widely speculated to be in the trillions of parameters [^2] GPT-4o and Gemini 1.5 Pro: How the New AI Models Compare - CNET. Notably, fine-tuning the smaller Llama-3.1-8B model on our benchmark’s training set resulted in substantial gains in both binary and free-answer tasks, with an average performance increase of 33.68% across the test set, even outperforming GPT-4o by 4.2%.
@@
Our study reveals a consistent pattern across all the LLMs examined, wherein their performance declines when tasked with questions involving negative fluents compared to those focused on fluents that are true. Specifically, we observed a 12.16% decrease in accuracy on these negative fluent tasks. Furthermore, when questions required reasoning about both true and false fluents simultaneously, LLMs exhibited competence in identifying the true fluents but demonstrated difficulty in correctly recalling the false ones.

### v3

@@ 1 Introduction
Table 1: Differences between ActionReasoningBench (Ours) and previous benchmarks on RAC. PlanBench (Valmeekam et al., 2024) ; TRAC (He et al., 2023)
@@ 2 Related Works
Evaluating the reasoning capabilities of LLMs using synthetic datasets has become a key focus in NLP, with increasing efforts to create challenging benchmarks. Notable areas of interest include the legal reasoning (Fei et al., 2023; Guha et al., 2023), logical reasoning (Luo et al., 2024; Han et al., 2024; Patel et al., 2024; Parmar et al., 2024) arithmetic reasoning (Cobbe et al., 2021; Miao et al., 2021), temporal reasoning (Uddin et al., 2024; Fatemi et al., 2024), and commonsense reasoning (Onoe et al., 2021; Lin et al., 2021; Geva et al., 2021; Lourie et al., 2021). Despite this progress, RAC remains significantly under-explored, even though it plays a crucial role in several of these reasoning tasks, such as commonsense and legal reasoning. To fill this gap, we create ActionReasoningBench using synthetically generated data.
Evaluating RAC and Planning
Benchmarking planning capabilities of LLMs is a well-studied area, with recent works (Zheng et al., 2024; Xie et al., 2024) demonstrating the challenges LLMs face. While planning is a non-polynomial problem that remains inherently difficult to solve, we believe that RAC, which is a polynomial problem, is a prerequisite for effective planning. Without comprehending the effects of actions, LLMs are unlikely to construct coherent plans. In this work, we address this gap by proposing ActionReasoningBench, a diagnostic benchmark for RAC.
Previous research, such as Banerjee et al. (2020), has investigated RAC capabilities in models like RoBERTa, focusing primarily on binary questions or single-word answers. Extending this, He et al. (2023) assessed LLMs on a broader range of question categories. However, only two of these—Action Executability and State Tracking—pertain directly to RAC, with the remainder addressing the broader domain of planning. Similarly, Valmeekam et al. (2024) introduced a benchmark evaluating both RAC and planning, with a detailed emphasis on planning tasks but maintaining a limited focus on RAC categories, specifically State Tracking and Action Executability. A comparative analysis of the benchmarks from He et al. (2023), Valmeekam et al. (2024), and our proposed benchmark is summarized in Table 1.
Parallel to our work, Kokel et al. (2024) introduced ACPBench to evaluate both RAC and planning. While ACPBench overlaps with our work, it primarily covers fundamental aspects of RAC. In contrast, our benchmark introduces greater complexity by incorporating categories such as State Tracking, Numerical RAC, Composite Questions, and constraints related to ramifications.
In this work, we concentrate on RAC and develop a benchmark that encompasses a wider range of categories, enabling more precise identification of areas where LLMs underperform. Furthermore, our benchmark introduces novel constraints involving ramifications, incorporating the indirect effects of actions. To the best of our knowledge, ramification constraints have not been addressed in any existing benchmark, marking a significant advancement in the evaluation of RAC capabilities.
@@ 3 ActionReasoningBench
This section provides a detailed overview of our benchmark, including its categorization, creation methodology, and validation process. A sample instance is presented in Appendix B, where we also describe the objects, actions, and fluents within the domain.
@@ 3.1 Question Categories
    State Tracking - This category extends the concept of Fluent Tracking. It involves querying about the complete set of fluents in the final state. For instance, in the Blocksworld domain, a state-tracking question might be “What are all the valid properties in this state?”
@@
    Action Executability - This category encompasses two types of questions related to executability of actions. The types of questions within this category are as follows:
@@ 3.2 Fluent Categories
Classification of every fluent across all 8 domains can be found in Appendix I. Furthermore, for each fluent type, we generate questions involving negative fluents, i.e. fluents that are false, which allows us to evaluate LLM’s comprehension of negation within RAC contexts.
@@ 3.3 Dataset Structure and Variations
ActionReasoningBench requires domains that facilitate the evaluation of LLMs on both short and long sequences of meaningful actions, where the effects and preconditions of these actions are succinctly described. Additionally, these domains should reflect real-world scenarios. To meet these criteria, we selected 8 domains Blocksworld, Depots, Driverlog, Grippers, Mystery, Satellite, Spanner, and Visitall–sourced from the International Planning Competition (IPC), covering the years 1998 to 2014. These domains are commonly used as benchmarks for evaluating advanced planning systems and provide a robust foundation for research in automated planning. Appendix G provides a detailed description of each domain. Notably, even state-of-the-art LLMs like GPT-4o are not capable of generating diverse domains or action sequences that conform to the precise constraints outlined in these domain descriptions, justifying the reliance on IPC domains.
@@
The domains provided by the IPC are described using the Planning Domain Definition Language (PDDL), a formal language designed to model deterministic actions and state transitions for planning problems. Further information on PDDL can be found in Appendix H.1. In this study, we initially translated the PDDL-based domains into natural language. Subsequently, ramification constraints were introduced into these natural language descriptions. The process underwent validation by two experts in the RAC domain to ensure correctness. Given that the category Action Executability focuses on determining whether the action can be performed rather than analyzing its effect, we concentrated on the categories Fluent Tracking, State Tracking, and Effects of Actions when introducing ramifications.
@@
We formulate two distinct types of questions based on the expected answer format. The first type consists of binary questions, where the response is either True or False. The second type involves subjective answers, which encompass a range of multiple objects, actions, or fluents.
@@ 3.4 Data Creation & Validation
The question generation process follows a four-stage pipeline, as illustrated in Figure 1. The selected domains from the IPC are represented in PDDL (see Appendix G for examples). First, these PDDL representations are used to generate 10 pairs of initial and goal conditions. A PDDL solver (Helmert, 2006) and validator (Howey et al., 2004) are then employed to obtain and validate the action sequences necessary to transition from the initial to the goal state. In the second stage, the PDDL domains, instances, and action sequences are converted into Answer Set Programming (ASP) descriptions using Python-based templates.
In the third stage, ASP solvers are used to generate the action-state space and extract fluents for each state, along with identifying all executable and inexecutable actions. Further details on these formal languages are provided in Appendix H. Finally, the fourth stage involves converting the action-state data into questions using a Python template. Up to three natural language variations are created for every object, action, and fluent. These templated sequences are then paraphrased to Llama-3.1-70B-Instruct to ensure they sound natural and avoid repetition in long action sequences. Three independent annotators review both the templated and the paraphrased versions to assess their naturalness, as detailed in Appendix D. Additionally, all eight domain descriptions were manually translated from PDDL to natural language.
@@ 3.5 Data Splits
Table 2: Overview of the test and train sets across Question and Fluent Categories. The “Mixed Fluents” category represents questions that involve more than one type of fluent.
@@ 4 Experiments and Evaluation
To evaluate our benchmark, we tested four LLMs and employed two prompting techniques. The LLMs include two proprietary models–GPT-4o (Achiam et al., 2023) and o1-preview (OpenAI, 2024)–alongside two open-source models, Llama-3.1-8B-Instruct and Llama-3.1-70B-Instruct (Dubey et al., 2024). Each LLM was evaluated using both few-shot prompting with three examples (few-shot-3) Brown (2020) and zero-shot-CoT Kojima et al. (2022) prompting.
While the entire dataset requires reasoning abilities, the ramification subset involves the most complex and challenging reasoning tasks. Given that o1-preview is specifically optimized for reasoning tasks and incurs significantly higher costs compared to GPT-4o [^1] As of Oct 2024, o1-preview is six times more expensive than GPT-4o https://openai.com/api/pricing/, we restricted its use to the ramification subset, where its superior reasoning capabilities are expected to provide the greatest benefit. Utilizing o1-preview across the entire dataset would not be cost-effective, as its advantages would be less pronounced for simpler reasoning tasks.
To ensure a fair evaluation of the standalone reasoning capabilities of LLMs, we deliberately avoided incorporating external tools or systems. Although integrating formal solvers (e.g., PDDL-based planners) with LLMs could potentially improve performance, our primary focus is on assessing the intrinsic reasoning abilities of these models.
@@
ActionReasoningBench includes two types of answer formats, as outlined in Section 3.3: binary (true/false) and free-form responses. The evaluation process was performed separately for each answer type. For binary questions, we extracted “true” and “false” keywords from the model’s response and compared them to the ground truth. Since free-form answers can’t be evaluated using exact string matching, we employed human evaluation for the ramification questions. While human evaluation is highly accurate, it is not scalable, so we used Llama-3.1-70B-Instruct to assess all free-form responses. The specific prompt used to evaluate the LLMs and the correlation between Llama-3.1-70B-Instruct and human evaluations are detailed in Appendix F TODO: Inter. For all experiments, we report the accuracy along with the standard error of the mean (SEM), calculated as $SEM=\frac{\sigma}{\sqrt{n}}$, where $\sigma$ represents the standard deviation, and $n$ is the sample size.
@@
We fine-tuned the Llama-3.1-8B model using the training data split outlined in Section 3.5. Due to the limited computing power, we excluded any data samples that exceeded a context length of 4096 tokens. The fine-tuning process was performed separately for free-response and binary questions. Detailed information on the fine-tuning procedure is provided in Appendix E. All experiments were executed using 8×H100 GPUs.
@@ 5 Results and Discussion
This section presents the results and analysis using ActionReasoningBench. The Zero-shot-CoT results for each LLM on both the binary and free-response subsets of the test set are provided in Tables 5 and 3 respectively. Similarly, the Few-shot-3 results for each LLM evaluated on these same subsets are displayed in Tables 6 and 7. The detailed analysis of the effects when ramification constraints are incorporated into the descriptions is discussed in Section 5.1.
@@
In our evaluation, GPT-4o demonstrated the highest performance on the Grippers domain and the lowest on the Satellite domain, with a performance gap of 15.53%. For both Llama-3.1-8B-Instruct and Llama-3.1-70B-Instruct, the best performance is also on the Grippers domain, but their worst performance occurs on the Mystery domain, with differences of 19.24% and 24.04%, respectively. Interestingly, the second lowest performing domain for GPT-4o is Mystery, while for the Llama models, it is Satellite. A detailed breakdown by domains is presented in Appendix C.3. This suggests that despite potential differences in pre-training data, these models exhibit a similar relative understanding of the domains.
Furthermore, when correlating these results with the State Space Complexity metric (described in App. G), traditionally used to gauge complexity for classical AI systems, we observe an inconsistent trend. This discrepancy implies that LLMs may employ different heuristics from those of traditional AI systems when tackling RAC problems, an observation that opens avenues for future exploration.
Figure 2: Performance for every Fluent Category for both binary and free-answer questions for every Action-Sequence length for GPT-4o, Llama-3.1-70B-Instruct, and Llama-3.1-8B-Instruct for Zero-shot-CoT prompt. Note: Bars represent SEM.
Performance across Question Categories
LLMs perform well in Fluent Tracking, State Tracking, and Effects of Actions, demonstrating their strength in keeping track of changes. However, they struggle with Action Executability, Composite Questions, and Numerical RAC. The performance of all LLMs on free-response complex questions highlights significant challenges, especially in the Numerical RAC category. This category reformulates existing question types into numerical formats, a domain where all tested LLMs exhibit notable difficulty. Specifically, performance on numerical questions related to Action Executability is 8.65% lower than on questions in the Fluent Tracking category. Previous research, such as (Ahn et al., 2024; McCoy et al., 2023), indicates that LLMs struggle with arithmetic reasoning and counting, which, when mixed with the RAC questions, likely contributes to the poor performance in the Numerical RAC category.
@@
As evident from Figure 2 and Table 9, across all LLMs examined in the study, a consistent trend emerges in which performance on Static Properties decreases as the length of action-sequence increases. While these static properties remain unchanged throughout the actions, they might get overlooked in longer sequences, likely due to their absence in the effect of any action. This phenomenon resembles the “needle in a haystack” challenge in long-context scenarios, where LLMs struggle to recall specific information embedded within a long context (Zhang et al., 2024). Conversely, Base Fluents maintain stable performance across all action sequences, indicating that the LLMs consistently capture the direct effects of actions. Ramification fluents exhibit a steady decline in performance as the sequence lengthens, particularly affecting the subcategory of Derived Fluents, which suggests that LLMs have more difficulty handling indirect effects. Finally, Mixed Fluents, which involve more than one fluent type, show a consistent decline in performance as the length of action sequences increases.
@@
Figure 3: Performance for every QUestion Category for both binary and free-answer questions for every Action-Sequence length for GPT-4o, Llama-3.1-70B-Instruct, and Llama-3.1-8B-Instruct for Zero-shot-CoT prompt. Note: Bars represent SEM.
Figure 3 illustrates the performance of the three models–GPT-4o, Llama-3.1-70B-Instruct, and Llama-3.1-8B-Instruct–across varying action sequence lengths in a zero-shot-CoT setting. The results combine both binary and free-response formats, with detailed data available in Tables 5 and 3. Generally, model accuracy declines as the action-sequence length increases, a pattern that holds for most categories. However, GPT-4o deviates from this trend in the State Tracking category, where performance first decreases and then improves. Since this trend is not observed with the other two models, and the results lie within the margin of error, we believe this is an outlier. In contrast, the Effects of Actions category consistently deviates from this trend, likely due to the nature of the task, which focuses on changes resulting from the last action, making it less dependent on the sequence of actions.
@@
As demonstrated in Tables 5 and 3, the Llama-3.1-70B-Instruct model consistently outperforms the smaller Llama-3.1-8B-Instruct model, with an average performance improvement of 20.84%. This improvement is likely due to the larger model’s superior reasoning capabilities stemming from its increased number of parameters. A similar trend is observed when comparing GPT-4o to Llama-3.1-70B-Instruct, where GPT-4o exhibits an average performance increase of 8.64%. Although the specific size of GPT-4o remains undisclosed, it is widely speculated to be in the trillions of parameters [^2] GPT-4o and Gemini 1.5 Pro: How the New AI Models Compare - CNET. Notably, fine-tuning the Llama-3.1-8B model on the training set resulted in substantial gains in both binary and free-answer tasks, with an average performance increase of 33.68% across the test set, even outperforming GPT-4o by 4.2%.
@@
Our study reveals a consistent pattern across all the LLMs examined as demonstrated in Table 10, wherein their performance declines when tasked with questions involving negative fluents compared to those focused on fluents that are true. Specifically, we observed a 12.16% decrease in accuracy on these negative fluent tasks. Furthermore, when questions required reasoning about both true and false fluents simultaneously, LLMs exhibited competence in identifying the true fluents but demonstrated difficulty in correctly recalling the false ones.